### v1

# Self-Ensembling Gaussian Splatting for Few-shot Novel View Synthesis

Abstract
3D Gaussian Splatting (3DGS) has demonstrated remarkable effectiveness for novel view synthesis (NVS). However, the 3DGS model tends to overfit when trained with sparse posed views, limiting its generalization capacity for broader pose variations. In this paper, we alleviate the overfitting problem by introducing a self-ensembling Gaussian Splatting (SE-GS) approach. We present two Gaussian Splatting models named the $\mathbf{\Sigma}$-model and the $\mathbf{\Delta}$-model. The $\mathbf{\Sigma}$-model serves as the primary model that generates novel-view images during inference. At the training stage, the $\mathbf{\Sigma}$-model is guided away from specific local optima by an uncertainty-aware perturbing strategy. We dynamically perturb the $\mathbf{\Delta}$-model based on the uncertainties of novel-view renderings across different training steps, resulting in diverse temporal models sampled from the Gaussian parameter space without additional training costs. The geometry of the $\mathbf{\Sigma}$-model is regularized by penalizing discrepancies between the $\mathbf{\Sigma}$-model and the temporal samples. Therefore, our SE-GS conducts an effective and efficient regularization across a large number of Gaussian Splatting models, resulting in a robust ensemble, the $\mathbf{\Sigma}$-model. Experimental results on the LLFF, Mip-NeRF360, DTU, and MVImgNet datasets show that our approach improves NVS quality with few-shot training views, outperforming existing state-of-the-art methods. The code is released at:https://github.com/sailor-z/SE-GS.
Figure 1: Qualitative results of our SE-GS and state-of-the-art approaches. The models are trained on sparse views and the images rendered from novel views are shown. As highlighted in the zoomed-in patches, our SE-GS captures finer details and produces fewer artifacts for novel views when trained on few-shot images.
1 Introduction
Novel view synthesis (NVS) is a critical task [49] in computer vision and graphics, playing a pivotal role in applications such as virtual reality [8], augmented reality [50], and 3D content generation [15, 42]. The objective of NVS is to generate photo-realistic images from previously unseen viewpoints. Typically, NVS starts by constructing a 3D representation [26, 39] from a set of existing 2D observations. In recent years, 3D Gaussian Splatting (3DGS) [20, 6, 47] has emerged as a powerful representation, integrating the advantages of both explicit [33] and implicit [26] representations. This approach enables efficient novel view generation and yields promising synthesized results with densely sampled observations that cover a wide range of viewpoints.
(a) Training
(b) Testing
Figure 2: Overfitting in 3D Gaussian Splatting with few-shot training views. 2(a) and 2(b) illustrate the performance of 3DGS on training and testing views, respectively. Each curve represents the PSNR values across different training iterations.
However, 3DGS tends to overfit the available views when only a limited number of images are provided. As shown in Fig. 2, we evaluate the 3DGS model trained on sparse images with different numbers of iterations. The performance on the training data consistently improves as the number of iterations increases, while the testing results deteriorate after 2000 iterations. Moreover, the overfitting issue becomes more noticeable with fewer training views, such as when using only 3 views. To address this problem, we introduce a new 3DGS approach in this paper, which enhances the quality of novel view synthesis with sparse training views via self-ensembling. Self-ensembling [11, 21, 27] has been witnessed in the literature as an effective strategy to mitigate overfitting, particularly in scenarios where limited data is available [45]. This technique leverages multiple predictions from the model under different conditions such as dropout [37] and data augmentation [36], forming a consensus prediction via regularization. By aggregating outputs with diverse perturbations, self-ensembling creates a more stable target for training, reducing the noise in individual predictions. Despite its success, how to utilize self-ensembling in the 3DGS pipeline remains an open problem. Therefore, we explore this gap, introducing a new 3DGS training pipeline consisting of a $\mathbf{\Sigma}$-model and a $\mathbf{\Delta}$-model.
Specifically, we present a $\mathbf{\Delta}$-model that represents a temporal sample in the Gaussian parameter space. Notably, in the scenario of sparse views, the training data lacks sufficient information to identify an optimal sample in the parameter space. As a result, the training process tends to converge to a local optimum, leading to an overfitting problem. To tackle this issue, one could create more samples by training multiple $\mathbf{\Delta}$-models, thereby increasing the likelihood of reaching a global optimum. Nevertheless, as we will demonstrate in Sec. 4, this approach incurs significant computational costs, limiting its ability to produce diverse samples. In contrast, we train a single $\mathbf{\Delta}$-model following the typical 3DGS training strategy [20], while obtaining various samples via perturbing the $\mathbf{\Delta}$-model. Rather than applying random perturbations, which often result in invalid samples, we perturb the $\mathbf{\Delta}$-model based on uncertainties derived from the training data. We store images generated by rendering the $\mathbf{\Delta}$-model from a set of pseudo views in buffers at different training iterations. These pseudo views are randomly sampled from the camera trajectories of the training views. We assess the reliability of the renderings by calculating the pixel-level uncertainties across the images within each buffer. Pixels with large uncertainty scores are identified, and the Gaussian parameters of the $\mathbf{\Delta}$-model associated with these pixels are perturbed by random noise. Therefore, our perturbation strategy generates uncertainty-aware temporal samples without introducing significant training overhead.
Building upon the $\mathbf{\Delta}$-model, we construct an ensemble of the temporal samples by proposing a $\mathbf{\Sigma}$-model. We train the $\mathbf{\Sigma}$-model over the training views without the aforementioned perturbations. To achieve self-ensembling, we regularize the geometry of the $\mathbf{\Sigma}$-model, minimizing the discrepancies between the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model. The discrepancies are measured through a photometric loss between images synthesized from the pseudo views. During testing, we employ the $\mathbf{\Sigma}$-model as the primary model for novel view synthesis. Since the regularization is performed based on diverse samples in the Gaussian parameter space, the self-ensembling process enhances the robustness of the $\mathbf{\Sigma}$-model, thereby improving the generalization ability to novel views.
We conduct experiments on several datasets consisting of LLFF [25], DTU [19], Mip-NeRF360 [3], and MVImgNet [46]. The evaluation is performed with varying numbers of training views. Our SE-GS achieves the best performance across all the datasets when limited training views are used, surpassing the state-of-the-art approaches. Additionally, we provide a comprehensive analysis in which the results demonstrate the effectiveness and efficiency of our method. To the best of our knowledge, we are the first to explore the potential of the ensembling mechanism in 3DGS for few-shot novel view synthesis. We present a new self-ensembling technique that effectively mitigates the overfitting of 3DGS without requiring additional ground-truth signals or incurring significant training overhead.
2 Related Work
Radiance field modeling. Recently, Neural Radiance Fields (NeRF) [26, 41, 30, 3] has brought groundbreaking advancements to novel view synthesis. NeRF learns the density and color information of a scene or object from multi-view posed images, enabling the synthesis of realistic images from novel views via volume rendering [9]. However, most NeRF-based methods are inefficient [17] during training and inference because they require numerous MLP queries in the rendering process. This limitation makes it challenging to support tasks with real-time requirements. To this end, Kebr et al. [20] proposes an efficient alternative for radiance field modeling, which enables neural rendering via 3D Gaussian Spaltting (3DGS). 3DGS relies on explicit representations that allow for faster rendering and training [10]. Some methods [40, 5, 24, 7] explore the generalization ability of 3DGS towards novel scenes, optimizing the Gaussian parameters based on learnable cost volumes [14]. In this paper, we stick to the vanilla 3DGS rendering pipeline, focusing on the scene-level radiance field modeling.
3DGS with few-shot images. The inadequacy of training views is a primary factor leading to artifacts in novel view synthesis. Even when training views are relatively dense, some regions may still be observed from few-shot images, resulting in quality degradation in the synthesized views. To handle such an under-constrained problem, some approaches incorporate additional ground-truth signals into the 3DGS pipeline. For instance, DNGaussian [23], CoherentGS [29], and FSGS [53] utilize a monocular depth estimator [32] to predict ground-truth depth maps, enabling the 3DGS model to be trained with both photometric and depth loss. Other methods, such as [43] and [16], leverage multi-view stereo techniques [35] to generate novel-view images as ground truth. However, ground-truth data obtained from off-the-shelf methods is inevitably noisy, which potentially impacts the training of 3DGS. As noted in [48] and observed in our experiments, when the number of training views increases, the performance of these methods is sometimes worse than the vanilla 3DGS. In this context, Zhang et al. [48] proposes training multiple 3DGS models with a cross-model regularization term. This method mitigates the overfitting problem without relying on additional data. Nevertheless, training additional 3DGS models leads to significant computational costs, limiting the effectiveness of the regularization.
Ensembling. Ensembling has been witnessed as a powerful technique [52, 1, 13] in machine learning to improve model robustness and generalization by aggregating predictions from multiple models. Traditionally, ensembles [12, 31, 22] are created by training independent models and averaging their outputs or applying a voting mechanism. To improve the efficiency, some self-ensembling methods, such as temporal ensembling [21] and consistency regularization [11], leverage variations of a single model across training iterations to build an ensemble. Motivated by the success of self-ensembling, we present a Self-Ensembling Gaussian Splatting approach, in which we generate diverse samples from the Gaussian parameter space through an uncertainty-aware perturbation mechanism.
3 Method
Figure 3: Pipeline of the presented SE-GS. We tackle the scenario where only sparse posed images are available. To mitigate the overfitting of 3DGS, we incorporate a self-ensembling mechanism into the 3DGS framework. We present a $\mathbf{\Delta}$-model that represents a temporal sample in the Gaussian parameter space. This $\mathbf{\Delta}$-model is perturbed during training, which results in diverse samples. The perturbation process is carried out with the guidance of uncertainties computed from images rendered from pseudo views. The pixels in the image buffers with high uncertainties are identified and the Gaussians overlapping these pixels are perturbed, highlighted as red ellipses. The self-ensembling over the temporal samples is achieved by training a $\mathbf{\Sigma}$-model with a regularization that penalizes the discrepancies of the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model. During inference, the novel view synthesis is performed using the $\mathbf{\Sigma}$-model.
3.1 Preliminaries and Motivations
3D Gaussian Splatting [20] represents a scene as a collection of Gaussians. These Gaussians are splatted onto the 2D image plane during rendering. Formally, we denote the set of $N$ Gaussians as $\{{G_{i}},i=1,2,...,N\}$. Each Gaussian $G_{i}$ is defined by $G_{i}=(\mathbf{\mu}_{i},\Sigma_{i},\mathbf{h}_{i},o_{i})$, where $\mathbf{\mu}_{i}$ is the center of the Gaussian in 3D space, $\Sigma_{i}$ is the covariance matrix, $\mathbf{h}_{i}$ represents spherical harmonics (SH) coefficients associated with the Gaussian, and $o_{i}$ indicates the opacity. Each Gaussian contributes to a 3D point $\mathbf{x}$ according to a 3D Gaussian distribution:
$G_{i}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mathbf{\mu}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{\mu}_{i})}.$ (1)
To ensure that $\Sigma_{i}$ remains positive semi-definite throughout optimization, this matrix is decomposed into two learnable components as $\Sigma_{i}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}$ stands for a scaling matrix. For RGB rendering, the projection from 3D space to a 2D image plane is defined as:
$\Sigma_{i}^{{}^{\prime}}=\mathbf{J}\mathbf{W}\Sigma_{i}\mathbf{W}^{T}\mathbf{J}^{T},$ (2)
where $\Sigma_{i}^{{}^{\prime}}$ is the projected 2D covariance matrix, $\mathbf{W}$ represents the projection matrix, and $\mathbf{J}$ denotes the Jacobian of the affine approximation of the projection. The color of a pixel is obtained by performing alpha-blending as
$\mathbf{c}=\sum_{i=1}^{M}\mathbf{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (3)
where $M$ is the number of Gaussians covering the pixel, $\mathbf{c}_{i}$ denotes the color of the Gaussian derived from the SH coefficients, and $\alpha_{i}$ is computed from the 2D covariance matrices and opacity scores. For better convergence, the Gaussian parameters are initialized based on the 3D points obtained by employing structure-from-motion (SFM) techniques [33, 34]. These parameters are optimized via gradient descent over the available posed training images.
In this paper, we focus on the scenario of sparse views, where only few-shot training images are known. We denote the 3DGS model trained on these images as $\mathcal{G}$, which represents a sample in the Gaussian parameter space. In this setting, $\mathcal{G}$ is prone to overfitting the training data and thereby getting trapped in local optima, which ultimately degrades the quality of the synthesized novel views.
To handle the overfitting problem, a promising solution [48] is to train a set of 3DGS models $\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{k}\}$ that correspond to different samples. By applying regularization across this set, each model is guided away from specific local optima, facilitating a more thorough exploration of the parameter space. As we will report in Sec. 4, the performance on testing data improves as $k$ grows. However, training multiple 3DGS models incurs significant computational costs, which become unaffordable with a large $k$. This limitation therefore acts as an obstacle to scaling up the method for better performance. Drawing inspiration from the success of the regularization, we propose a novel training strategy based on self-ensembling [11, 21], which enhances 3DGS for few-shot NVS with negligible additional training overhead. The pipeline is illustrated in Fig. 3 and the details will be elaborated in this section.
3.2 Temporal Samples in 3DGS
Figure 4: Buffer update during training. For each sampled pseudo view, we dynamically update the buffer which stores the images rendered at different training steps. For instance, at training step $t_{T}$, the oldest image $\mathbf{I}_{T-S}$ in the buffer is popped, and the new image $\mathbf{I}_{T}$ is pushed into the buffer. An uncertainty map $\mathbf{U}^{t_{T}}$ is computed based on the current buffer, which is then employed for the perturbation.
In contrast to generating multiple samples by training separate 3DGS models, we dynamically create temporal samples from a single $\mathbf{\Delta}$-model during training. Specifically, we train the $\mathbf{\Delta}$-model on the available posed images, following the same optimization and density control strategies introduced in 3DGS [20]. At each training iteration, the current $\mathbf{\Delta}$-model represents a specific sample in the Gaussian parameter space based on the information contained in the training data. A temporal 3DGS model is then sampled from the neighboring region of the $\mathbf{\Delta}$-model in the parameter space by perturbing the $\mathbf{\Delta}$-model as:
$\hat{G}_{\Delta}^{t}=G_{\Delta}^{t}+\delta_{t},$ (4)
where $G_{\Delta}^{t}$ stands for a Gaussian in the $\mathbf{\Delta}$-model at the training step $t$, $\delta_{t}$ indicates the perturbations, and $\hat{G}_{\Delta}^{t}$ denotes a perturbed Gaussian in the temporal model. To achieve this, a plausible approach is to sample random noise from a normal distribution, $\delta_{t}\in\mathcal{N}(\mu_{t},\sigma_{t})$. However, since these perturbations are independent of the training data, this method tends to result in invalid samples.
Consequently, we present an uncertainty-aware perturbing strategy according to the statistics on the renderings. The approach starts by sampling pseudo views along the camera trajectories of the training views. Given two cameras sampled from the training views, the camera extrinsic of a pseudo view is computed as:
$\hat{\mathbf{R}}=\text{SLERP}(\mathbf{R}_{1},\mathbf{R}_{2},\beta),$ (5)
$\hat{\mathbf{c}}=\beta\mathbf{c}_{1}+(1-\beta)\mathbf{c}_{2},$ (6)
$\hat{\mathbf{T}}=-\hat{\mathbf{R}}\hat{\mathbf{c}},$ (7)
where $(\mathbf{R}_{1},\mathbf{R}_{2})$ represent the rotations of the sampled cameras, $(\mathbf{c}_{1},\mathbf{c}_{2})$ indicate the sampled camera centers, SLERP denotes the spherical linear interpolation [4], $\beta$ is a scalar that controls the interpolation, and $(\hat{\mathbf{R}},\hat{\mathbf{T}})$ stand for the camera parameters of the pseudo view. As illustrated in Fig. 3, for each pseudo view, we render an RGB image using the current $\mathbf{\Delta}$-model and store the renderings at different training steps in a buffer. The uncertainties of these renderings are then computed as:
$\mathbf{U}=\frac{1}{S}\sqrt{\sum_{i=1}^{S}(\mathbf{I}_{i}-\bar{\mathbf{I}})^{2}},$ (8)
where $\mathbf{U}\in\mathbb{R}^{H\times W}$ is the uncertainty map with the same resolution as the rendered image, $\mathbf{I}_{i}$ represents an image in the buffer, $\bar{\mathbf{I}}$ indicates the mean of these images, and $S$ is the buffer size. This uncertainty estimation is carried out over all sampled pseudo views in parallel. In practice, to enhance the robustness, we perform local smoothing over each uncertainty map, which is denoted as:
$\hat{\mathbf{U}}(i,j)=\frac{1}{k^{2}}\sum_{m,n}\mathbf{U}(i+m,j+n),$ (9)
where $k=5$ is the kernel size. Subsequently, we apply the uncertainty-aware perturbations to each of the Gaussians in the $\mathbf{\Delta}$-model, formulated as:
$\hat{G}_{\Delta}^{t}=G_{\Delta}^{t}+\delta_{t}h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t}),$ (10)
where $\hat{\mathcal{U}}^{t}$ is a set of the computed uncertainty maps at the current training step, and $h(\cdot,\cdot)$ is an indicator function defined as:
$h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t})=\begin{cases}1&\text{if }\max\limits_{(i,j)\in\mathcal{P}(G_{\Delta}^{t})}u_{ij}\geq\tau\\ 0&\text{if }\max\limits_{(i,j)\in\mathcal{P}(G_{\Delta}^{t})}u_{ij}<\tau,\end{cases}$ (11)
in which $\tau$ is a predefined threshold and $u_{ij}$ is the uncertainty score of a pixel sampled from a set of pixels $\mathcal{P}$ overlapping with the 2D splat of $G_{\Delta}^{t}$. Please refer to the supplementary material for details on defining $\tau$ and identifying $\mathcal{P}$. Therefore, the reliability of the Gaussians is connected with the uncertainties of renderings, and only the unreliable Gaussians characterized by high uncertainty scores are perturbed. In practice, we add the perturbations to the 3D points, rotations, scales, and opacities of the $\mathbf{\Delta}$-model. Particularly, since the 3D rotation is not contiguous when formulated as a 3D matrix, we perturb the 6D continuous representation [51] which is defined as:
$\hat{\mathbf{R}}^{t}_{\Delta}=f^{-1}(f(\mathbf{R}_{\Delta}^{t})+\delta^{R}_{t}h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t})),\ \ \delta^{R}_{t}\in\mathbb{R}^{6}.$ (12)
Fig. 4 illustrates the dynamic updates of the buffer throughout the training process. As the buffer is updated, the uncertainty map varies accordingly, resulting in diverse temporal samples from the $\mathbf{\Delta}$-model.
Moreover, it is worth noting that the number of Gaussians in the $\mathbf{\Delta}$-model varies during training due to density control, making it challenging to directly compute the uncertainties of the Gaussian parameters across different training steps. The proposed perturbing approach can handle this issue since the statistics on 2D renderings provide a consistent and efficient assessment, regardless of the varying numbers of Gaussians.
3.3 Self-Ensembling in 3DGS
Given the temporal samples around the $\mathbf{\Delta}$-model, the next step in our pipeline is to construct an ensemble from these samples. Concretely, we train a separate 3DGS model named the $\mathbf{\Sigma}$-model on the posed images without perturbations. To prevent the $\mathbf{\Sigma}$-model from being stuck at local optima, we guide the training of the $\mathbf{\Sigma}$-model by leveraging a regularization technique [48]. The regularization is formulated as:
$\mathcal{L}_{r}=(1-\lambda){||\mathbf{I}^{t}_{\Sigma}-\mathbf{I}^{t}_{\Delta}||}_{1}+\lambda\mathcal{L}_{\text{D-SSIM}}(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta}),$ (13)
where $\lambda=0.2$ is a predefined weight, $\mathcal{L}_{\text{D-SSIM}}$ denotes a D-SSIM term, and $(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta})$ represent the images rendered from a pseudo view using the current $\mathbf{\Sigma}$-model and $\mathbf{\Delta}$-model, respectively. We also utilize a co-pruning strategy [48] to enhance the regularization. Therefore, the $\mathbf{\Sigma}$-model aggregates information from different samples in the parameter space, serving as an ensemble of 3DGS models. The final loss function during training is defined as:
$\mathcal{L}=\mathcal{L}_{\text{RGB}}+\gamma\mathcal{L}_{r},$ (14)
where $\gamma=1$ by default and $\mathcal{L}_{\text{RGB}}$ represents a photometric loss over the training views. At testing, we keep the $\mathbf{\Sigma}$-model for novel view synthesis. Notably, in contrast to CoR-GS, we derive these samples from a single $\mathbf{\Delta}$-model without additional training costs. As a result, our SE-GS can be efficiently scaled up to a large number of 3DGS models, enabling more effective regularization. Moreover, our regularization is carried out on pseudo views in a self-supervised manner, which is thus independent of additional ground truth such as monocular depth [23, 53].
[…]
3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
Mip-NeRF [2] 16.11 22.91 24.88 0.401 0.756 0.826 0.460 0.213 0.160 0.215 0.090 0.066
DietNeRF [18] 14.94 21.75 24.28 0.370 0.717 0.801 0.496 0.248 0.183 0.240 0.105 0.073
RegNeRF [28] 19.08 23.10 24.86 0.587 0.760 0.820 0.336 0.206 0.161 0.149 0.086 0.067
FreeNeRF [44] 19.63 23.73 25.13 0.612 0.779 0.827 0.308 0.195 0.160 0.134 0.075 0.064
SparseNeRF [38] 19.86 - - 0.624 - - 0.328 - - 0.127 - -
3DGS [20] 19.22 23.80 25.44 0.649 0.814 0.860 0.229 0.125 0.096 0.120 0.066 0.051
DNGaussian [23] 19.12 - - 0.591 - - 0.294 - - 0.132 - -
FSGS [53] 20.43 24.09 25.31 0.682 0.823 0.860 0.248 0.145 0.122 0.104 0.066 0.054
CoR-GS [48] 20.45 24.49 26.06 0.712 0.837 0.874 0.196 0.115 0.089 0.101 0.060 0.046
SE-GS 20.79 24.78 26.36 0.724 0.839 0.878 0.183 0.110 0.084 0.100 0.059 0.045
Table 1: Results on LLFF with 3, 6, and 9 training views. We highlight the best, second-best, and third-best results in red, orange, and yellow, respectively.
[…]
3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
3DGS [20] 17.67 23.69 26.80 0.804 0.894 0.941 0.158 0.086 0.050 0.112 0.058 0.031
MVSplat [7] 17.33 16.34 16.22 0.598 0.596 0.587 0.279 0.296 0.304 - - -
DNGaussian [23] 18.57 - - 0.776 - - 0.178 - - 0.110 - -
FSGS [53] 17.84 23.68 26.17 0.822 0.905 0.941 0.161 0.096 0.064 0.110 0.056 0.036
CoR-GS [48] 18.65 24.39 27.38 0.835 0.910 0.950 0.140 0.074 0.045 0.099 0.049 0.028
SE-GS 19.24 25.28 28.08 0.857 0.924 0.958 0.132 0.073 0.043 0.089 0.045 0.026
Table 2: Results on DTU with 3, 6, and 9 training views. We use red, orange, and yellow to indicate the best, second-best, and third-best results, respectively. Object masks are used for all evaluated methods to remove background when conducting the evaluation.
4 Experiments
4.1 Setup
Datasets. We conduct experiments on four datasets, i.e., LLFF [25], DTU [19], Mip-NeRF360 [3], and MVImgNet [46]. On LLFF, DTU, and Mip-NeRF360, we follow the experimental setup introduced in CoR-GS [48], using the same training/testing splits and downsampling rates. As suggested in CoR-GS, we mask the background when assessing the quality of novel view synthesis on the DTU dataset. Since LLFF, DTU, and Mip-NeRF360 offer a limited number of scenarios, we extend our experiment to a large-scale dataset, i.e., MVImgNet. We randomly sample 50 scenes from this dataset and resize the longest side of each image to 512. The Gaussian parameters for all evaluated methods are initialized based on the point clouds obtained from COLMAP [34] to ensure a fair comparison. It is worth noting that COLMAP fails on certain scenes within the DTU dataset in the setting of sparse views, where we instead randomly initialize the point clouds. Please refer to the supplementary material for more details on the setup.
Implementation details. We train our SE-GS for 10,000 iterations on the LLFF, DTU, and MVImgNet datasets, and for 30,000 iterations on Mip-NeRF360. In the uncertainty-aware perturbing mechanism, we employ 24 image buffers with a buffer size of 3 and perturb the $\mathbf{\Delta}$-model with an interval of 500 iterations. The noise $\delta_{t}$ in Eq. 10 is sampled from a standard normal distribution, i.e., $\delta_{t}\in\mathcal{N}(0,1)$. More details are provided in the supplementary material.
Method 12-view 24-view
[…]
3DGS [20] 18.52 0.523 0.415 0.167 22.80 0.708 0.276 0.096
FSGS [53] 18.80 0.531 0.418 0.163 23.28 0.715 0.274 0.093
CoR-GS [48] 19.52 0.558 0.418 0.151 23.39 0.727 0.271 0.091
SE-GS 19.91 0.596 0.400 0.143 23.74 0.745 0.265 0.086
Table 3: Results on Mip-NeRF360 with 12 and 24 training views. The first, second, and third-best results are marked in red, orange, and yellow, respectively.
[…]
5-view 7-view 10-view 5-view 7-view 10-view 5-view 7-view 10-view
3DGS [20] 26.48 29.06 31.82 0.819 0.878 0.924 0.301 0.228 0.184
FSGS [53] 26.88 29.12 31.68 0.848 0.886 0.922 0.358 0.292 0.250
CoR-GS [48] 27.62 29.77 32.09 0.842 0.889 0.928 0.302 0.235 0.190
SE-GS 27.88 30.19 32.68 0.857 0.902 0.937 0.277 0.214 0.177
Table 4: Results on MVImgNet with 5, 7, and 10 training views. Results colored in red, orange, and yellow denote the best, second-best, and third-best performances, respectively.
4.2 Quantitative results
We list the quantitative results of the evaluated approaches on the LLFF, DTU, Mip-NeRF360, and MVImgNet datasets in Table 1, Table 2, Table 3, and Table 4, respectively. The best, second-best, and third-best results in each column are highlighted in red, orange, and yellow, respectively. For MVSplat [7], we employ the pretrained model released by the authors and evaluate it on DTU. Notably, the per-scene optimized 3DGS methods significantly outperform MVSplat, highlighting the advantages of scene-level radiance field modeling for high-quality novel view synthesis. The performance of both NeRF and 3DGS methods consistently declines as the number of training views decreases, underscoring the challenge of few-shot novel view synthesis. In this context, our method surpasses all competitors across the four datasets. Note that in the scenario of sparse views, the available information in the training data is limited, making it more challenging to improve performance compared to the dense-view setting.
Specifically, on the LLFF and DTU datasets, we train the methods using 3, 6, and 9 views. Our method achieves an improvement of $0.34$ in terms of PSNR with 3 views on LLFF. The improvement is more significant on DTU, reaching a gain of $0.59$. Since the images in the Mip-NeRF360 dataset are captured from diverse viewpoints, we sample more training views, i.e., 12 and 24, following the setting used in CoR-GS. On this dataset, our SE-GS achieves the highest PSNR and SSIM, as well as the lowest LPIPS and AVGE, demonstrating superior NVS quality under the challenges of sparse views with large-scale viewpoint variations. Additionally, to further assess the robustness across diverse scenarios, we conduct an experiment on 50 scenes sampled from the MVImgNet dataset. As reported in Table 4, our method obtains consistently better results than previous state-of-the-art approaches when trained with 5, 7, and 10 views. Our SE-GS also demonstrates better compatibility with relatively dense views. For instance, the PSNR improvement increases from $0.26$ to $0.59$ as the number of training views varies from 5 to 10.
Methods like DNGaussian [23] and FSGS [53], which leverage auxiliary data terms sometimes outperform 3DGS. However, in some cases, such as 9 views on the LLFF dataset, the PSNR of FSGS becomes worse than 3DGS. Since the data acquired through off-the-shelf approaches might be unreliable, the quality of the NVS results is consequently affected. Compared with FSGS and DNGaussian, CoR-GS [48] exhibits better stability, outperforming 3DGS in more scenarios. This finding highlights the promising potential of regularization in the sparse-view setting. However, the improvement is still not consistent, as CoR-GS shows worse LPIPS than 3DGS when trained with 10 views on the MVImgNet dataset. In contrast, our SE-GS beats 3DGS in all cases, demonstrating superior stability. Our method is capable of avoiding specific local optima since the $\mathbf{\Sigma}$-model aggregates information from diverse temporal samples in the presented self-ensembling mechanism.
4.3 Qualitative results.
Figure 5: Qualitative results. The methods are trained on sparse views and the renderings of novel views are illustrated. The images are selected from the DTU and MVImgNet datasets.
Fig. 5 presents qualitative comparisons among 3DGS, CoR-GS, and our SE-GS, illustrating renderings from novel views. The models are trained on the DTU and MVImgNet datasets with 3 and 5 posed images, respectively. Our SE-GS produces fewer visual artifacts compared to the other methods, achieving better robustness against the challenge of few-shot training views. Moreover, our method captures finer details, particularly in areas with complex and repeated textures such as the pineapple and watch strap. This demonstrates the effectiveness of our self-ensembling strategy in improving both the visual quality and stability of novel view neural rendering under limited-view conditions.
(a) PSNR
(b) Training Speed
Figure 6: Comparison with CoR-GS. CoR-GS [48] is trained with different numbers of Gaussian models. The number varies from 2 to 8. PSNR on LLFF with 3 training views is utilized as the metric to evaluate the NVS quality. Iterations/s refers to the number of training iterations completed per second, indicating the training speed.
4.4 Analysis
To shed more light on the effectiveness and efficiency of our SE-GS, we develop a comprehensive analysis of the introduced self-ensembling mechanism. The analysis starts by comparing our approach with CoR-GS trained with varying numbers of Gaussian models. The detailed results on LLFF with 3 training views are shown in Fig. 6. As illustrated in Fig. 6(a), the PNSR of CoR-GS increases as more Gaussian models are trained. Since adding more Gaussian models enhances the regularization effect in CoR-GS, this observation highlights the potential of improving 3DGS in the sparse-view setting by strengthening the regularization process. The PSNR reaches the peak with 6 Gaussian models, indicating an upper bound when using a limited number of models. In contrast, our method perturbs the $\mathbf{\Delta}$-model based on the uncertainties computed from dynamically updated image buffers, which results in much more diverse samples of Gaussian models compared to CoR-GS. This leads to more effective regularization, as reflected by the better PSNR score.
Moreover, Fig. 6(b) shows the training speed measured by the number of training iterations completed per second. As witnessed in this figure, the speed of CoR-GS significantly drops as the number of Gaussian models increases. This limitation poses a challenge in scaling CoR-GS up to a large number of Gaussian models. Note that in our pipeline, only two models are trained, i.e., the $\mathbf{\Delta}$-model and $\mathbf{\Sigma}$-model. The temporal samples are generated through perturbations. The result in Fig. 6(b) shows that we achieve a comparable training speed to CoR-GS with 2 models. Only a negligible additional cost is incurred due to the perturbation process. It is worth noting that only the $\mathbf{\Sigma}$-model is retained for neural rendering during inference, so the testing speed of SE-GS is the same as that of the original 3DGS. Consequently, our method is capable of efficiently enhancing the performance of 3DGS with sparse training views.
Method Random Gradient-Aware Uncertainty-Aware
[…]
Table 5: Effectiveness of the uncertainty-aware perturbations. Random means we add random perturbations to all Gaussians in the $\mathbf{\Delta}$-model. Gradient-Aware indicates an alternative in which we perturb the Gaussians based on gradients. All methods are trained on LLFF with 3 views for 2000 iterations and the PSNR is utilized as the metric.
Furthermore, we assess the effectiveness of our uncertainty-aware perturbing strategy, comparing it with two alternatives. Specifically, as we introduced in Sec. 3, the straightforward method perturbing the $\mathbf{\Delta}$-model is to add random noise to all Gaussians in the model. We denote this approach as Random. Building upon this baseline, an alternative approach, referred to as Gradient-Aware, identifies unreliable Gaussians by analyzing their gradients. Specifically, we replace the indicator function in Eq. 11 with one based on gradient magnitude. Gaussians with gradients exceeding a threshold are perturbed. In practice, we adopt the same threshold as used in density control. We train these two alternatives and our approaches on LLFF with 3 views for 2000 iterations, which challenges the methods in terms of convergence speed. As listed in Table 5, Random performs the worst since it leads to many invalid Gaussian samples, negatively impacting convergence speed. Note that our uncertainty-aware strategy aggregates information from multiple training steps, while the gradient-aware method can only utilize the gradients at the current training step. The improvement in PSNR shown in Table 5 demonstrates the superior effectiveness of our approach.
5 Conclusion
In this paper, we have introduced a new self-ensembling mechanism that enhances the novel view synthesis of 3DGS with few-shot training images. We tackle the challenge of overfitting by training an ensemble model named $\mathbf{\Sigma}$-model with the guidance of diverse temporal samples from the Gaussian parameter space. We introduce an effective sampling strategy that perturbs a $\mathbf{\Delta}$-model based on the uncertainties computed from dynamically updated buffers of pseudo-view renderings. We have conducted experiments on LLFF, DTU, Mip-NeRF360, and MVImgNet, as well as a comprehensive analysis of our approach. The experimental results have demonstrated the effectiveness, stability, and efficiency of our SE-GS.
Acknowledgment. This work was funded in part by the Swiss National Science Foundation via the Sinergia grant CRSII5-180359 and the Swiss Innovation Agency (Innosuisse) via the BRIDGE Discovery grant 40B2-0_194729.